Solve the equation.
14) 4(2z - 3) = 7(z + 5)

Answer by edjones(8007)

4(2z - 3) = 7(z + 5)
8z-12=7z+35
z=47 Subtract 7z and add 12 to each side.
Check:
4(2*47-3)=7(47+5)...4(94-3)=7*52...4*91=364...364=364
Ed